Neverstop Corporation sells item A as part of its product line. Information about the beginning inventory, purchases, and sal4. Prepare journal entries to record the purchase and sale transactions, as well as the cost of sales, assuming that the weigAssume that because of a clerical error, the ending inventory is reported to be 505 units rather than the actual number of unI just need the answers to 5a. and 5b.
Answer #1

5a. Number of units reported in ending inventory are less

Value of Ending Inventory at 505 units= 365*4.05 + 140*3.90

= $2,024.25

Value at 605 units = 365*4.05 + 240*3.90 = $2,414.25

Hence, ending Inventory is understated by 2414.25 – 2024.25 = $390

Hence, cost of sales will be Overstated by $390

5b. Current Assets i.e. ending Inventory will be understated by $390